i need help in finding the common solution for these two systems 2x-3y=15 and x+y=6 ?
Found 2 solutions by mananth, sudhanshu_kmr:
Answer by mananth(16949)

2x-3y=15
x+y=6
multiply by 3
3x+3y=18
add this equation to the first equation
3x+2x+3y-3y=15+18
5x=33
x= 33/5
..
plug value og x in second equation
..
33/5 +y = 6
33+5y / 5 =6
33+5y = 30
5y=30-33
5y=-3
/5
y = -3/5
..
m.ananth@hotmail.ca

Answer by sudhanshu_kmr(1152)


multiply second equation by 3 and add to first

2x -3y = 15
3x + 3y = 18 ( multiply by 3)
now after addition 5x = 33 => x = 33/5
now put the value of x on any equation of above ,
for example in x+y = 6
=> 33/5 +y =6
=> y = 6 - 33/5
=> y = ( 30-33)/5
=> y = -3/5

thus, x = 33/5, y = -3/5